OpenStudy (jjuden):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 13, c = 21

OpenStudy (jjuden):

a 15.83 b 15.87 c 16.19 d 16.49

OpenStudy (anonymous):

a^2+b^2=c^2 so... a^2+13^2=21^2 a^2+169=441 a^2=441-169 a^2=272 a=sqrt(272) a=16.49

OpenStudy (mathmale):

Please invest some effort into finding a solution. You might draw this triangle, along with the known side lengths, and then determine how you'd find the missing side.

OpenStudy (anonymous):

 \[\sqrt{13^2+21^2}\]

OpenStudy (anonymous):

That will = A

OpenStudy (jjuden):

@nohegarcia thank you for showing the steps i kinda understand now